A train left Town A at 8:50 a.m., it passed Town B at 9:30 a.m. and arrived at Town C at 10:30 a.m. The distance between Town A and Town B is 100 km. Between Town B and Town C, the train travelled at 110 km/h. What was the average speed of the train for the whole journey?

Answer by mananth(12270)

A to B is 100 km
B to c speed 110 km/h traveled for 1 hour so distance = 110 km.
Total distance = 210 km
Total time = 10:30 - 8:50 = 1 h 40 minutes = 1 2/3 hours===> 5/3 hours
Rate = distance / time
Rate = 210/(5/3)
Rate = 210*3/5
Rate = 42*3
Rate = 126 km/h ====> average speed of the journey